An investigation of 23713 households was made in an urban and rural mixed locality. Of these 1618 were farmers, 2015 well to do and 770 families were having at least one graduate. Of these graduate families 335 were those of farmers and 428 were well to do; also 587 well to do families were those of farmers and out of them only 156 were having at least one graduate. Obtain all the ultimate class frequencies.


1
Expert's answer
2022-06-10T11:16:53-0400
"P(F)=\dfrac{1618}{23713}\approx0.0682"

"P(W)=\dfrac{2015}{23713}\approx0.0850"

"P(G)=\dfrac{770}{23713}\approx0.0325"

"P(F|G)=\dfrac{335}{770}\approx0.4351"

"P(W|G)=\dfrac{428}{770}\approx0.5558"

"P(F|W)=\dfrac{587}{1618}\approx0.3628"

"P(G|W)=\dfrac{156}{1618}\approx0.0964"